# End-to-End Predictive Planner for Autonomous Driving with Consistency Models

### v1

@@ Abstract
Trajectory prediction and planning are fundamental components for autonomous vehicles to navigate safely and efficiently in dynamic environments. Traditionally, these components have often been treated as separate modules, limiting the ability to perform interactive planning and leading to computational inefficiency in multi-agent scenarios. In this paper, we present a novel unified and data-driven framework that integrates prediction and planning with a single consistency model. Trained on real-world human driving datasets, our consistency model generates samples from high-dimensional, multimodal joint trajectory distributions of the ego and multiple surrounding agents, enabling end-to-end predictive planning. It effectively produces interactive behaviors, such as proactive nudging and yielding to ensure both safe and efficient interactions with other road users. To incorporate additional planning constraints on the ego vehicle, we propose an alternating direction method for multi-objective guidance in online guided sampling. Compared to diffusion models, our consistency model achieves better performance with fewer sampling steps, making it more suitable for real-time deployment. Experimental results on Waymo Open Motion Dataset (WOMD) demonstrate our method’s superiority in trajectory quality, constraint satisfaction, and interactive behavior compared to various existing approaches.
@@ I Introduction
Figure 1: The overview of our method.
@@
To navigate safely and efficiently in a dynamic environment, autonomous vehicles must effectively predict and interact with diverse road participants, including other vehicles, pedestrians, and cyclists. This typically requires a prediction module to anticipate other agents’ future trajectories and a planning module to generate trajectories for the ego vehicle. While prediction modules have widely adopted data-driven approaches to learn from human-driving data, planning modules typically rely on optimization-based methods. These two modules usually operate in a decoupled, alternating manner [31, 10].
This decoupled method introduces fundamental limitations. The interactive behavior generated by the planner is inherently reactive rather than proactive, as the planner responds to other agents without accounting for how they might react to the ego vehicle. This limitation can lead to the computationally expensive theory of mind (TOM) reasoning [53, 12, 17, 18]. In highly interactive scenarios, such as lane merging where proactive nudging is essential, reactive planning may cause the ego vehicle to be stuck [3, 5]. Existing solutions either struggle to scale with the number of agents [25] or compromise optimality due to a restrictive exploration of the solution search space [49].
In addition, modular approaches optimize prediction and planning independently, which often leads to less effective behavior compared to jointly trained, end-to-end frameworks [30]. The alternation between prediction and planning modules also introduces latency and computational inefficiency, making it challenging to meet the demands of real-time operation.
Our key insight is that modeling the joint distribution of future trajectories for both the ego vehicle and surrounding agents can unify prediction and planning into a single, end-to-end framework. In this approach, the ego trajectory, informed by the desired goal location, represents the plan, while the trajectories of surrounding agents serve as predictions. Additional planning constraints can be imposed on the ego trajectory as needed. When trained on large-scale human driving datasets like the Waymo Open Motion Dataset (WOMD) [15], this joint distribution naturally captures the complex interactive behaviors observed in real-world scenarios.
Diffusion models [51, 20] offer a promising method to sample from this high-dimensional, multimodal joint distribution of ego and surrounding agents’ trajectories. They have demonstrated success in generating high-quality images and videos [39, 21] or robot trajectories [38, 27]. Moreover, their ability to model conditional distributions makes them well-suited for trajectory planning applications, as it allows for the incorporation of crucial context for planning and prediction, including trajectory history, map information, ego vehicle’s goal locations, etc. Compared to transformers, diffusion-based method also supports controllable generation to accommodate additional requirements solely at test time through guided sampling, without any additional training.
@@
However, diffusion models typically require many sampling steps to obtain high-quality samples, making it challenging to meet the demands of real-time operations on autonomous vehicles. To overcome this limitation, consistency models [52] are designed to generate data directly from noise in a single step or a small number of steps, significantly improving sampling efficiency without compromising quality.
In this paper, we present an end-to-end predictive planner based on consistency models, designed to unify planning and prediction within a single data-driven framework. Our method samples from the conditional joint distribution of ego and surrounding agents’ future trajectories, given encoded map and trajectory history along with the ego vehicle’s goal configuration. While we use a transformer-based scene encoder [45] in our implementation, our approach is encoder-agnostic and can integrate with any suitable scene encoder. Trained on WOMD [15], our method ensures proactive trajectory planning by capturing complex interactions between agents observed in the real world. Additionally, during online testing, planning constraints are incorporated through guided sampling using our proposed alternating direction method for multi-objective guidance.
Experiments demonstrate that our consistency model-based planner generates effective interactive behaviors with other agents, such as nudging and yielding, through a single inference, eliminating the need for iterative alternation between prediction and planning modules. It outperforms diffusion models by requiring fewer sampling steps while delivering superior trajectory quality. Compared to transformer-based models, it enables controllable generation through online guided sampling that significantly improves planning constraint satisfaction.
@@ II Related Work
For trajectory predictions, data-driven methods have shown promising results when trained on large-scale real-world human driving datasets such as Waymo Open Motion Dataset (WOMD) [15], INTERACTION Dataset [62], Argoverse [8] or nuScenes [7]. Graph Neural Networks have been used to model the semantic information in traffic scenarios for interactive prediction [23]. Generative models like Variational Autoencoders and Diffusion Models have also been utilized for multimodal trajectory prediction tasks [43, 27]. Recently, transformers [56], particularly Motion Transformers (MTR) [45, 46, 34] with encoder-decoder architectures, have achieved state-of-the-art results in trajectory prediction [58].
@@
In trajectory planning, the goal is to determine a feasible and efficient path for the ego vehicle, based on observations of the environment and surrounding agents. This task is often formulated as a numerical optimization problem, typically decomposed path planning and speed planning to facilitate real-time operation in dynamic environments [60, 61, 2]. Game-theoretic approaches have been explored to model agent interactions [41], with efficient computation [16] that considers imperfect traffic agents [54], safe actions [55], and uncertainty reduction in the estimates of agent behaviors [22].
Beyond standalone planning, preliminary effort has been made to integrate trajectory predictions into planner design for collision avoidance [31] and proactive motion planning [4, 44, 49]. While the aforementioned methods have advanced trajectory prediction and planning performance in dense environments, to our best knowledge, none have attempted to unify these two to achieve human-like interactive behavior with computational efficiency. Instead, our work addresses this gap by presenting a data-driven framework for end-to-end joint prediction and planning with a single inference.
@@
Diffusion models [47, 51, 20] are generative models capable of producing high-quality samples from complex distributions, initially achieving success in image and video generations [39, 21]. The Denoising Diffusion Probabilistic Model (DDPM) [20] gradually corrupts the training data with Gaussian noise and learns to reverse this process through iterative denoising with deep neural networks. These models commonly use U-Net [40] or transformer-based architectures [56] as backbones. Conditional sampling in diffusion models is enabled by techniques such as classifier guidance [13] and classifier-free guidance [19] However, DDPM requires numerous sampling steps to preserve the sample quality, resulting in significant computational costs. To address this inefficiency, Denoising Diffusion Implicit Models (DDIM) [48] introduces a deterministic generation process that balances sampling quality and speed, while progressive distillation [42] reduces the number of sampling steps by distilling a pre-trained diffusion model into a more efficient one. Consistency model [52, 50] takes a different approach, directly generating data from noise. They can be trained through distillation of a pre-trained diffusion model or a standalone consistency training and outperform distillation methods in one- and few-step sampling [52].
Diffusion and consistency models also have shown success in robotics and autonomous systems for their expressiveness and ability to provide samples from complex distributions. In robotics, conditional diffusion models have been applied to policy learning [26, 1, 11], with consistency models following similar applications [14, 37]. Diffusion models have also been adopted for trajectory optimization across various robotics tasks [33, 32]. In autonomous driving, diffusion models have shown promising results in trajectory prediction [27, 57] and controllable traffic simulation or scenario generation [63, 59, 9, 24], capturing complex interactive behaviors in traffic scenarios. However, our work differs fundamentally from these efforts: we focus on developing a predictive trajectory planner with consistency models, leveraging guided sampling to incorporate planning constraints. Unlike scenario generation tasks, where sampling efficiency is less critical, our consistency model-based predictive planner needs significantly shorter sampling steps to deliver high-quality interactive behaviors compared to diffusion models.
@@ III Problem Formulation
We aim to plan the trajectory of the ego vehicle with a planning goal, simultaneously predicting the interactive behaviors of others. We assume the trajectory histories of traffic agents and map information are given and known.
At the current timestep $t_{0}\in\mathbb{R}$, the ego agent’s state history over the past $H_{1}\in\mathbb{N}$ timesteps including the current state is given as $S_{e}^{\text{hist}}=(s_{e}^{t_{0}},s_{e}^{t_{0}-1},...,s_{e}^{t_{0}-H_{1}})\in\mathbb{R}^{(H_{1}+1)\times d_{s}}$, where $d_{s}$ is the dimension of the states, e.g. position, velocity, etc. Let $\mathcal{O}$ denote the set of other agents of all types such as vehicles, pedestrians, and cyclists. Their history states are given for the current timestep $t_{0}$ and the past $H_{1}$ timesteps, denoted by $S_{\mathcal{O}}^{\text{hist}}=(s_{o}^{t_{0}},s_{o}^{t_{0}-1},...,s_{o}^{t_{0}-H_{1}})_{o\in\mathcal{O}}\in\mathbb{R}^{N_{\mathcal{O}}\times(H_{1}+1)\times d_{s}}$, where $N_{\mathcal{O}}=|\mathcal{O}|$. The map information is denoted by $M\in\mathbb{R}^{N_{m}\times c\times d_{m}}$, which includes $N_{m}\in\mathbb{N}$ polylines in the scenario with $c$ points on each polyline and $d_{m}$ attributes per point. We use fixed dimensions $N_{\mathcal{O}}$ and $N_{m}$ for agents and map polylines respectively, which are chosen large enough to accommodate all scenarios with padding and masks applied to invalid entries. We assume the ego agent’s goal position is given by some high-level planner and is notated $s_{e}^{g}\in\mathbb{R}^{d_{s}}$.
@@ IV Preliminaries
IV-A Motion-Transformer Encoder
To encode the trajectory history $(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})$ and map information $M$ as the conditional input for the consistency model, we adopt the encoder architecture from MTR [45, 46]. This transformer-based architecture effectively models scene context through a local attention mechanism, enabling efficient interaction modeling between agents and the road map while maintaining memory efficiency. It also introduces a dense prediction head with a loss function $\mathcal{L}_{\text{encoder}}$ to train this encoder individually. For more detailed descriptions of this encoder, we refer readers to the original MTR paper [45]. The MTR++ encoder [46] would be suitable for multi-agent settings, but its implementation is not yet available. More importantly, our method is encoder-agnostic, allowing for flexible choice of any suitable scene encoder to integrate with our consistency model-based predictive planner. We represent the encoded feature of the trajectory history and map with MTR encoder as $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})\in\mathbb{R}^{(N_{\mathcal{O}}+1)\times D}$ and $\text{MTR}(M)\in\mathbb{R}^{N_{m}\times D}$, respectively, where $D$ is the embedded feature dimension.
@@ IV-B Consistency Model
The consistency model can generate high-quality samples from complex distributions with only one or a few sampling steps [52]. It consists of a forward diffusion process and a reverse process. Let $\mathcal{D}$ be our dataset of trajectories. We first normalize all trajectories in $\mathcal{D}$ using the empirical mean and standard deviation computed across the entire dataset. Let $\mathcal{X}$ be the space of such normalized trajectories, and $p_{\text{data}}$ be the data distribution over $\mathcal{X}$. In the forward process, we first draw an initial sample $x_{1}$ from $p_{\text{data}}$. We then apply an increasing noise schedule ${\sigma_{1},\sigma_{2},...,\sigma_{T}}$ with $\sigma_{1}\approx 0$ to progressively corrupt $x_{1}$ through $T$ steps. Specifically, at each step $i$, we draw a noise sample $\epsilon$ from $\mathcal{N}(0,I)$ and set the corrupted data $x_{i}$ as:
$\quad x_{i}=x_{1}+\sigma_{i}\cdot\epsilon,\quad i=1,2,...,T$ (1)
We choose $\sigma_{T}$ large enough so that, when $x_{1}$ is repeatedly sampled from $p_{\text{data}}$ and corrupted via additive Gaussian noise, the empirical distribution of the resulting $x_{T}$ approximates $\mathcal{N}(0,\sigma_{T}^{2}I)$.
Let $\mathcal{Y}$ be the space of conditioning information. In the reverse process, we aim to learn a consistency function $f_{\theta}$ with the parameter $\theta$ that maps a noisy trajectory sample $x_{i}$, condition $y\in\mathcal{Y}$, and the noise level $\sigma_{i}$ directly to the corresponding clean sample $x_{1}$ for each step $i=1,..,T$. This is achieved by choosing a specific function form of $f_{\theta}$:
$f_{\theta}(x,y,\sigma)=c_{\text{skip}}(\sigma)x+c_{\text{out}}(\sigma)F_{\theta}(x,y,\sigma)$ (2)
where $\sigma$ represents the noise level, and the differentiable function $c_{\text{skip}}$ and $c_{\text{out}}$ are chosen such that $c_{\text{skip}}(\sigma_{1})=1,c_{\text{out}}(\sigma_{1})=0$ when $\sigma=\sigma_{1}$ [28]. This ensures that the boundary condition of $f_{\theta}$ is met: $f_{\theta}(x_{1},y,\sigma_{1})=x_{1}$. $F_{\theta}$ is usually approximated by the neural network using a U-Net [40] or a transformer [36] based architecture.
For the consistency model training, we aim to enforce the consistency of the output from $f_{\theta}$ for adjacent sampling steps $i$ and $i+1$ to be both $x_{1}$. Let $\mathcal{P}[1,T-1]$ be a distribution over the integer set ${1,2,...,T-1}$. The consistency training minimizes the following loss function $\mathcal{L}_{\text{consistency}}$:
@@
where $\mathcal{P}$ can be the uniform distribution or lognormal distribution, and $d$ is chosen to be a pseudo-huber metric function $d(x,y)=\sqrt{||x-y||_{2}^{2}+x^{2}}-c$ [50].
During the data generation process, we first draw a sample $x_{T}$ from $\mathcal{N}(0,\sigma_{T}^{2}I)$. Then with the trained consistency model $f_{\theta}$, we perform iterative sampling by first predicting the approximate clean data $x_{1}^{[i]}$ and then sample $x_{i-1}$ with $\epsilon$ drawn from $\mathcal{N}(0,I)$:
@@
for $i=T,T-1,...,1$, until we obtain the final clean sample $x_{1}^{[1]}$.

### v2

# Predictive Planner for Autonomous Driving with Consistency Models

@@ Abstract
Trajectory prediction and planning are essential for autonomous vehicles to navigate safely and efficiently in dynamic environments. Traditional approaches often treat them separately, limiting the ability for interactive planning. While recent diffusion-based generative models have shown promise in multi-agent trajectory generation, their slow sampling is less suitable for high-frequency planning tasks. In this paper, we leverage the consistency model to build a predictive planner that samples from a joint distribution of ego and surrounding agents, conditioned on the ego vehicle’s navigational goal. Trained on real-world human driving datasets, our consistency model generates higher-quality trajectories with fewer sampling steps than standard diffusion models, making it more suitable for real-time deployment. To enforce multiple planning constraints simultaneously on the ego trajectory, a novel online guided sampling approach inspired by the Alternating Direction Method of Multipliers (ADMM) is introduced. Evaluated on the Waymo Open Motion Dataset (WOMD), our method enables proactive behavior such as nudging and yielding, and also demonstrates smoother, safer, and more efficient trajectories and satisfaction of multiple constraints under a limited computational budget.
@@ I Introduction
Figure 1: Our consistency model takes the map and agents history encoded with the planning goal, and generates a trajectory plan for the ego vehicle and predictions for surroundings. Multiple constraints are enforced through online guided sampling.
@@
To navigate safely and efficiently in a dynamic environment, autonomous vehicles must effectively predict and interact with diverse road participants, including other vehicles, pedestrians, and cyclists. This typically requires a prediction module to anticipate other agents’ future trajectories and a planning module to generate trajectories for the ego vehicle.
Traditionally, these two modules usually operate in a decoupled, alternating manner [27, 9], which introduces fundamental limitations. Some approaches integrate data-driven predictions into an optimization-based planner that is inherently reactive rather than proactive [15]. In highly interactive scenarios, such as lane merging where proactive nudging is essential, reactive planning may cause the ego vehicle to be stuck [4, 5]. Also, existing work attempts to incorporate the ego vehicle’s plan into the prediction module [18], but the initial ego plans may already be suboptimal if ignoring future interaction with others. In addition, the alternation between prediction and planning introduces computational inefficiency, making it challenging to meet the demands of real-time operation.
Recently, there has been growing interest in multi-agent trajectory generation using purely data-driven approaches. Transformer-based methods have been widely adopted for multi-agent prediction tasks [35, 28]. Diffusion models, with their strong ability to sample from high-dimensional multimodal distributions, have been adopted that improve the prediction performance [22]. Compared to transformer-based models, diffusion models also support controllable generation through guidance methods, making them particularly suitable for scenario generation [19, 47, 8].
@@
However, diffusion models typically require many sampling steps to obtain high-quality samples, which becomes challenging to meet the demands of real-time operations on autonomous vehicles. To address this limitation, consistency models [40] are designed to generate data directly from noise in a single step or a small number of steps. This significantly improves sample efficiency and maintains high sample quality under limited computational budgets[12, 29].
In this paper, we present a predictive planner for autonomous driving with a unified data-driven framework. Our key idea is to leverage a consistency model to sample the joint distribution of the ego and surrounding agents’ future trajectories, given the map, trajectory history, and the ego vehicle’s goal. Consistency models enable high-quality trajectory generation with only a few sampling steps, making them more suitable for online planning than standard diffusion models. Trained on WOMD [13], our method captures complex agent interactions and produces proactive trajectories observed in real-world driving. To further enforce multiple planning constraints, we introduce a novel online guided sampling approach inspired by ADMM [6]. Our planner is also encoder-agnostic and can be integrated with any suitable scene encoder.
Experiments show that our consistency model-based planner generates proactive behaviors, e.g., nudging and yielding, through a single forward pass without alternation between prediction and planning. It outperforms both diffusion models and an accelerated variant DDIM [37] with smoother, safer, and more efficient trajectories under limited computational budgets. Compared to transformer-based models, we enable controllable generation that significantly improves satisfaction of multiple trajectory constraints without affecting the original distribution.
@@ II Related Work
Traditionally, trajectory planning is often formulated as a numerical optimization problem, typically decomposed into path planning and speed planning to facilitate real-time operation in dynamic environments [2]. Beyond standalone planning, preliminary effort have integrated trajectory predictions into planner design for collision avoidance [27] and proactive motion planning [3, 33, 38], but either struggle to scale with the number of agents [20] or compromise optimality due to a restrictive exploration of the solution search space [38]. Game-theoretic approaches have been explored to model agent interactions [31], with efficient computation [14] that considers imperfect traffic agents [43], safe actions [42], and uncertainty reduction in the estimates of agent behaviors [17].
Recently, data-driven methods have shown promising results in trajectory predictions when trained on large-scale real-world human driving datasets such as Waymo Open Motion Dataset (WOMD) [13], nuScenes [7], etc. Transformer [44] based methods, such as Scene Transformer [28] and Motion Transformers (MTR) [34], use an encoder-decoder architecture and achieved great performances on Waymo Motion Prediction Challenge [45]. They allow for multi-agent trajectory generation [28, 35] with potential for goal-conditioned trajectory planning. However, trajectory planning usually requires more controllable generation to satisfy constraints provided online, which remains challenging for transformer-based approaches.
Diffusion models are powerful generative models capable of producing high-quality samples from complex distributions, initially gaining success in computer vision [36, 41, 16]. The Denoising Diffusion Probabilistic Model (DDPM) [16] corrupts training data with Gaussian noise and learns to reverse this process through iterative denoising. Controllable generation in diffusion models can be achieved through classifier guidance [11] during sampling. As a result, Diffusion models have been applied to robot policy learning [21, 1, 10] and trajectory generation [26, 25, 46]. In autonomous driving, diffusion model facilitates multimodal trajectory predictions [22] and controllable traffic scenario generation [47, 8, 19], capturing complex interactive behaviors.
@@
However, diffusion models typically require a large number of sampling steps to maintain sample quality, resulting in substantial computational overhead. To address this inefficiency, Denoising Diffusion Implicit Models (DDIM) [37] introduces a deterministic sampling process that balances speed and quality, while progressive distillation [32] reduces the number of sampling steps by distilling a pre-trained diffusion model into a more efficient one. Consistency model [40, 39] takes a different approach by directly generating data from noise in a single or a few steps. They outperform distillation methods [40] in image generation and have shown early success in robot policy learning [12, 29].
In this work, we focus on building a goal-conditioned predictive planner using a consistency model. Unlike offline scenario generation, where sampling efficiency is less critical, our method produces high-quality interactive behaviors with significantly fewer sampling steps than diffusion models, making it well-suited for high-frequency online planning. Compared to diffusion-based predictors, which model the full multimodal future distribution, our approach focuses on goal-conditioned generation tailored for planning. In contrast to transformer-based models, which lack controllable generation at inference time, our approach enables multiple planning constraints, e.g., goal-reaching and control limits, to be satisfied simultaneously through an ADMM-inspired online guided sampling.
@@ III Problem Formulation
Assuming an ego vehicle’s goal is given from a high-level planner and the trajectory histories of all traffic agents and map information are known, we aim to plan the trajectory of the ego vehicle while simultaneously predicting the interactive behaviors of others.
At the current timestep $t_{0}\in\mathbb{R}$, the ego agent’s state history over the past $H_{1}\in\mathbb{N}$ timesteps including the current state is given as $S_{e}^{\text{hist}}=(s_{e}^{t_{0}},s_{e}^{t_{0}-1},...,s_{e}^{t_{0}-H_{1}})\in\mathbb{R}^{(H_{1}+1)\times d_{s}}$, where $d_{s}$ is the dimension of the states, e.g. position, velocity, etc. Let $\mathcal{O}$ denote the set of other agents of all types, such as vehicles, pedestrians, and cyclists. Their history states are given for the current timestep $t_{0}$ and the past $H_{1}$ timesteps, denoted by $S_{\mathcal{O}}^{\text{hist}}=(s_{o}^{t_{0}},s_{o}^{t_{0}-1},...,s_{o}^{t_{0}-H_{1}})_{o\in\mathcal{O}}\in\mathbb{R}^{N_{\mathcal{O}}\times(H_{1}+1)\times d_{s}}$, where $N_{\mathcal{O}}=|\mathcal{O}|$. The map information is denoted by $M\in\mathbb{R}^{N_{m}\times c\times d_{m}}$, which includes $N_{m}\in\mathbb{N}$ polylines in the scenario with $c$ points on each polyline and $d_{m}$ attributes per point. We pad and mask the invalid entries. We assume the ego agent’s goal position is given as $s_{e}^{g}\in\mathbb{R}^{d_{s}}$.
@@ IV Preliminaries
IV-A Motion Transformer Encoder
To encode the trajectory history $(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})$ and map information $M$ as the conditional input for the consistency model, we adopt the encoder architecture from MTR [34, 35]. This transformer-based encoder effectively models scene context and agents’ interactions, and introduces a dense prediction head with a loss function $\mathcal{L}_{\text{encoder}}$ to train this encoder individually. The MTR++ encoder [35] would be suitable for multi-agent settings, but its implementation is not yet available. More importantly, our method is encoder-agnostic, allowing for flexible choice of any suitable scene encoder to be integrated. We represent this encoded feature of the trajectory history and map as $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})\in\mathbb{R}^{(N_{\mathcal{O}}+1)\times D}$ and $\text{MTR}(M)\in\mathbb{R}^{N_{m}\times D}$, respectively, where $D$ is the embedded feature dimension.
@@ IV-B Consistency Model
The consistency model can generate high-quality samples from complex distributions with only one or a few sampling steps [40]. It consists of a forward diffusion process and a reverse process. Let $p_{\text{data}}$ be the data distribution with zero mean and unit variance over space $\mathcal{X}$. In the forward process, we first draw an initial sample $x_{1}$ from $p_{\text{data}}$, then gradually corrupt it with an increasing noise schedule ${\sigma_{1},\sigma_{2},...,\sigma_{T}}$ through $T$ steps, with $\sigma_{1}\approx 0$. At each step $i$, the corrupted data is $x_{i}=x_{1}+\sigma_{i}\cdot\epsilon$ with $\epsilon$ sampled from $\mathcal{N}(0,I)$. $\sigma_{T}$ is chosen large enough such that the empirical distribution of the resulting $x_{T}$ approximates $\mathcal{N}(0,\sigma_{T}^{2}I)$.
Let $\mathcal{Y}$ be the space of conditioning information. In the reverse process, we aim to learn a consistency function $f_{\theta}$ with the parameter $\theta$ that maps a noisy trajectory sample $x_{i}$, condition $y\in\mathcal{Y}$, and the noise level $\sigma_{i}$ directly to the corresponding clean sample $x_{1}$ for each step $i=1,..,T$. This is achieved by choosing a specific function form of $f_{\theta}\coloneqq c_{\text{skip}}(\sigma)x+c_{\text{out}}(\sigma)F_{\theta}(x,y,\sigma)$, where $c_{\text{skip}}$ and $c_{\text{out}}$ are differentiable function with $c_{\text{skip}}(\sigma_{1})=1,c_{\text{out}}(\sigma_{1})=0$ [23], such that the boundary condition is met: $f_{\theta}(x_{1},y,\sigma_{1})=x_{1}$. $F_{\theta}$ is usually approximated by neural networks.
For the consistency model training, we aim to enforce the consistency of the output from $f_{\theta}$ for adjacent sampling steps $i$ and $i+1$ to be both $x_{1}$, through the following loss $\mathcal{L}_{\text{consistency}}$:
@@
where $\mathcal{P}[1,T-1]$ is lognormal distribution over the integer set $\{1,2,...,T-1\}$, and $d$ is a pseudo-huber metric function $d(x,y)=\sqrt{||x-y||_{2}^{2}+x^{2}}-c$ [39].
For data generation, we draw a sample $x_{T}$ from $\mathcal{N}(0,\sigma_{T}^{2}I)$. Then with the trained consistency model $f_{\theta}$, we perform iterative sampling by first predicting the approximate clean data $x_{1}^{[i]}$ and then sample $x_{i-1}$ with $\epsilon$ drawn from $\mathcal{N}(0,I)$:
@@
for $i=T,T-1,...,1$, until we obtain the clean sample $x_{1}^{[1]}$.